convert 22.27 degrees to degrees, minutes and seconds. round to the nearest second.
Answer by lwsshak3(11628)

convert 22.27 degrees to degrees, minutes and seconds. round to the nearest second.
.27 deg=.27*3600=972"=16.2'=16'12"
22.27 degrees=22˚16'12"
note:
1 min=60 sec
1 deg=60 min
1 deg=60*60=3600 sec